HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

LINE SHAPE CALCULATIONS FOR THE H_{2}O-O_{2} COLLISION SYSTEM

\underline{R. GAMACHE} \underline{N. ORPHANOS} , University of Massachusetts Lowell, 265 Riverside Street, 01854, Lowell, MA, USA; \underline{B. VISPOEL} , Research unit Lasers and Spectroscopies (LLS), Institute of Life, Earth and Environment (ILEE), University of Namur, 61 Rue de Bruxelles, Namur, Belgium; \underline{K. SUNG}, \underline{G. TOON}, Jet Propulsion Laboratory, California Institute of Technology, 4800 Oak Grove Drive, Pasadena,

CA 91109-8099, USA

Modified Complex Robert-Bonamy (MCRB) calculations were made for the H_{2}O-O_{2} collision system using an intermolecular potential comprised of dipole-quadrupole, quadrupole-quadrupole, atom-atom component expanded to 20^{th} order and rank 4, induction and London dispersion terms. The intermolecular potential was refined using the high-precision measurements of half-widths and line shifts made in this work. The final potential gives results that compare very well with the measurements; average percent difference of 0.052 and a standard deviation of 5.108 percent. With previous data sets the agreement was never better than ∼15 percent standard deviation. Calculations employing this intermolecular potential were then made for some fourteen thousand water vapor transition in the 5680-9090 cm^{−1} range and for the temperature range 200-3000 K for the H_{2}O-O_{2} collision systems. These data were combined with H_{2}O-N_{2} data to produce line shape data for H_{2}O-air. The temperature dependence of the half-widths and line shifts were determined with the Double Power Law model of Gamache and Vispoel^{1}. The final line shape data were added to a HITRAN-like line list for use in modeling tellurics features toward extreme precision radial velocity measurements made by the Palomar Radial Velocity Instrument (PARVI), which is a near infrared fiber-fed diffraction limited high resolution spectrograph built for Doppler measurements of the host stars of extrasolar planets.

1doi:10.1016/j.jqsrt.2018.05.019, R.R. Gamache and B. Vispoel, J. Quant. Spectrosc. Radiat.

Transfer 217, 440-452, (2018).

p-number: p029  Submitted on Sat, 03 Jun 2023 16:53:05 +0200